Find an equation in slope-intercept form passing through the points (2,5) and (7,-3)
Any help will be appreciated.

Answer by scott8148(6628)

first, find the slope __ y change (rise) divided by x change (run) __ (-3-5)/(7-2) __ -8/5

write in point-slope form __ y-5=(-8/5)(x-2)

solve for y (slope-intercept) __ y-5=(-8/5)x+(16/5) __ y=(-8/5)x+(41/5)